# Creative Writers’ Attitudes on Writing as Training Data for Large Language Models

### v1

@@ 5.1.2. Respect
While writers understood that by publishing their work they are allowing others to use that work in perhaps unexpected ways, they also felt that the use of their work as training data indicated a lack of respect both for writing as an endeavor and writers as people. Although writers did not express that all uses of their writing require consent—for instance, as in the section above about the creative chain, they expect that others will learn from their work or perhaps even try to mimic it—the use of their writing as training data felt different. P6 said that the use of their writing for AI may be more like signing away their foreign subrights, which would require consent and be part of contract negotiations when publishing.
Mechanisms of respect: compensation is key for professionalized writers. The idea that writers wanted respect was very much tied to the mechanisms in which respect can be shown. The mechanisms that writers desired was a function of many things: their degree of professionalization, the type of writing they worked on, their sense of precarity in the labor market, and their perception of the organizations that were or could be using their writing. Primarily, writers wanted respect for writing as a laborious act on its own—to write a book takes time, effort, and expertise. As P9 put it,
@@
The people we talked to who were professional writers—who earned the majority of the income from their creative writing practice—were most likely to cite compensation as the main mechanism of respect they would expect. For instance, P8, a professional freelance writer, noted that, “And I say this as a, you know, a very pro AI person who thinks it’s the next creative way. But content producers actually need to be compensated on some level.” P11 made the comparison with other kinds of expertise: “If NASA wanted my expertise to hire me to help in the space program they would actually have to pay me quite a lot of money, and they would do so.” However, technology companies were not treating writers as if they brought any kind of expertise or as if their labor was valuable.
@@
Power differentials modulate desire for compensation. Writers were also quick to point out the large power differential between the corporations creating these models and the writers trying to earn a living. For instance, P7 explains that, “The tech companies are making money hand over fist, and I feel like they can compensate people.” Professional writers were likely to say they would require compensation regardless of whether or not the model creator was making money, but many writers did make a distinction between organizations they perceived could afford to compensate writers, and non-profit organizations or applications. We return to the idea of who can afford to compensate writers in the realistic expectations section.
@@
Different kinds of writing require different kinds of respect. Writers discussed genre distinctions in terms of the level of labor involved. P9 talked about the use of certain kinds of books feeling particularly disturbing, “like Pulitzer Prize-winning work that’s happening in nonfiction,” which represents decades of research then being taken without consent or compensation and made “part of the churn.” In this case, writers were talking about respect for time-consuming labor that it represented.
@@
Other writers noted that the use of certain kinds of writing felt more disrespectful than others because of the contents of the writing, rather than the labor taken to produce it. P4 talked about some writing being “deeply personal” and being “disturbed by the lack of care” that is shown by the act of putting such work into a language model. P32, in discussing a colleague who had written a nonfiction book which included accounts of assault, noted that their feelings and words about the person who assaulted them now lives on in the predictive models trained on that book. ”There is an aspect of that that is like literal haunting” (P32). Here, respect was not about compensation but rather about understanding the sensitive nature of some kinds of writing.
@@
How to value writing. Writers discussed how they might consider whether or not to contribute certain works of theirs, which led them to express how different kinds of writing may be valued differently. P14 said that if they were asked for consent, “there are things that I would probably hold back in order to wait until I’ve gotten the value out of them and then at that point I wouldn’t worry about it as much.” Similarly P13 noted that they used to post a lot more of their ideas on social media before they started winning awards for their short stories; now, they consider their ideas to be more valuable and are more hesitant about sharing them publicly. Other writers noted that older writing is often less valuable, either, as P14 noted, because they have already gotten value out of it, or because older writing is worse—they have become a better writer and therefore value their more recent writing more (P7).
@@ 5.1.3. The Human Element
Many writers see the text generation of LLMs as an attempt to mimic human processes of writing, and, in many cases, an author’s unique style or voice. This is tied to discussions of the creative chain, as noted above, and writers view the training of LLMs on a particular writer’s voice as impoverished and decontextualized. When they read writing that inspires them, writers appreciate, contextualize, question, compare, and transform a text within their own social, embodied, and emotional contexts, something mass aggregates of textual data cannot do. P6 contrasts these processes: “My absolute favorite book was instrumental in the way that I wrote the favorite of all of the books I’ve written. But it’s still coming through me. Whereas, I feel like ChatGPT would end up poorly mimicking me, as opposed to being something that came from someone else, but was formed by their appreciation for me.”
@@
Implicit in these distinctions is an understanding that LLMs do not have authorial intent, which is central to writing as a human activity. That LLMs do not have authorial intent when generating text is viewed as an asset to those writers choosing to use them in creative ways. P31 notes this explicitly, while problematizing how LLMs are colloquially talked about as “collaborators” in some contexts. They say:
@@
The audience for a training dataset of creative work is not the audience that writers want or imagine. Undergirding our conversations with writers was the understanding that the imagined or intended audience for a writer’s work is other people, not a dataset using writing to train an LLM. Every direct mention of the words ‘audience’ or ‘reader’ in our interviews was in reference to an explicitly human audience and at least one of these two terms was mentioned in all of the interviews we conducted.
In light of the strong connection between humanity and writers’ conceptions of audience, writers were left to conjecture about who the ‘audience’ of training data really was. Algorithms, corporations, and users of LLMs were all variously floated as often ideas of what might constitute the readership for training datasets. P26 wants to share their work with people, but is “not ready to put [even] a little piece of it into the algorithm.” The scale and combinatory nature of training datasets provoked skepticism and sometimes fear around this imagined ‘readership,’ quite unlike the readerships writers imagine for themselves.
@@ 5.2.1. Lack of Control
Certain uses of writing are restricted and there are avenues for recourse. Writers often sign contracts that outline a publisher’s right to republish a book in a foreign country, or create a movie adaption, indicating that writers do have control over a variety of use cases of their writing. For instance, P11 said, “There are lots of sources of pirated books on the Internet, but they’re illegal, and we try to shut them down. If ever I find a free copy of my book circulated somewhere, I’ll ask my publisher to send a cease and desist letter, and it usually gets taken down.
@@
It is difficult to understand how to restrict only certain use cases of LLMs. Writers did not always disagree with the creation of language models per se, and instead had concerns only about certain use cases. But many writers noted that it seems impossible to prevent certain use cases. For instance, P24 would consider consenting to the use of their writing if they could ensure that the model would not be used to compete with them, but felt “like that would be an impossible ask. People are just gonna lie and go ‘No, I’m not writing a novel.’” Others noted that while it may be possible to audit big companies, “there are all these people putting out small models, anyone can do this now, and it almost feels impossible to rein in” (P2).
@@
This is what tech companies have always done. Uber is like, we’re just gonna make these cars. And we’ll see if cities can catch up with us. Or Google Maps when they’re like, we’re just gonna photograph everybody’s house. And then we’ll see if anyone gets mad. I think that’s totally what AI writing software is doing right. They’re waiting to see what people are gonna do. (P12)
@@
Finally, writers acknowledged the difficulty of predicting future use cases. As technology changes, it can be hard to know what you might be agreeing to in the moment. Writers were especially concerned about those at the start of their careers—established writers were concerned about up-and-coming authors (P20, P9) and writers we interviewed who were at the start of their careers noted their particular predicament (P13). As an early stage writer, one has less power to negotiate certain terms out of their contract and may end up agreeing to terms that, years down the line, perhaps when the writer is more famous, are unexpected, unprecedented, and ultimately undesirable.
5.2.2. Career Impacts
@@
Present day impacts of language models on the creative writing industry. Though public discourse about the impact of language models on the creative writing industry often takes place in the future, writers described present day impacts of language models on their careers and fields. P12 described the strain language models have placed on literary editors, who “spend hours every week” sifting through large amounts of “junk pitches and generated stories.” P12 recalled how this influx of generated work led one prominent literary magazine (Clarkesworld) to “close submissions entirely,” which “really hurts people who are just getting started, people who already have this narrowing slate of options where they can submit their work.” Other writers expressed similar concern over the impacts of language models on new writers getting started in the industry (P9).
@@
Writers with academic teaching positions described confronting the impacts of language models on creative writing education. These writers explained how their creative writing careers and teaching careers are inextricable, either because they are hired to teach because of their creative writing careers, or because they see teaching as part of their creative practice. These writers note that even if they are morally against LLMs, for instance because of the energy used to train and use them, they must deal with this technology in their classrooms (P32, P16, P14, P12, P18).
@@
Some writers also described positive impacts of language models and AI on their careers. Self-published authors described using language models and other forms of generative AI to help with the work a traditional publisher might do, like creating a cover or producing marketing materials for their work (P27, P24). P31, who makes use of language models in their work, described how their career has benefited from advances in LLMs and the resulting increase in public interest in AI.
@@
Predicted future impacts. By and large, most participants found that in their current state, language models are not direct competition for writers, as they do not produce sufficiently sophisticated, high quality writing. When speculating about how future advanced versions of language models might impact their careers and industries, many writers imagined competing with a language model that had been trained to mimic their writing. Other writers speculated more broadly about how language models might compete with or displace writers, i.e. P12: ”an AI that writes better novels than humans.
@@
I don’t think [LLMs are] particularly useful for the creative industry. I mean, I could see them being useful for, like, … discovering a new kind of a new antibiotic or something like that… But, human expression? I guess I think of poetry or fiction as a deeply human activity. And it sort of baffles me why people would want to read what has been historically a human activity that has been produced by a machine. (P10)
@@
Backdrop of a changing industry. Some writers acknowledged that their resistance to contribute to training datasets was “compounded by a sense of precarity” in the creative writing industry (P3). P20 noted that “it’s a really difficult time for people in our profession and for creative industries in general. The job market is constantly shrinking.
Some writers expressed protectiveness towards the creative industry, citing the value of humanity intrinsic to creative work. The questions these writers were asking were existential, implying that the existence of AI-generated art and literature is forcing us to ask, ‘What kind of world do we want to live in?’ P9 notes:
@@
P13, along with several others we spoke with, wants to live in a world where people can make a living as writers: “I don’t think it’s good to make it cheaper to make art for the fact that people need to be employed and I would prefer to be employed as a writer than as a lawyer.
@@ 5.2.3. Interpretation of Scale
Compensation is hard to determine and likely to be insignificant given the scale of LLMs. While most writers objected to their data being collected by corporations, they simultaneously many had difficulty comprehending how to value their contribution when considering the scale of training data used. Some noted that their style was not distinct enough to be noticeable in any outputs or to significantly change the model (P2, P24). P10 described this as, “It’s hard for me to trace how money being made is based on the training data.” Others felt it was hard to feel personally slighted when their contribution is so very small (P29). Still others were unsure exactly how compensation would even work: “It’s not like a spotify thing where you can be like, Okay, they listen to this track, or they read this chapter” (P17).
Despite many writers wanting compensation for the use of their work, they acknowledged that such compensation is likely to be small given the amount of data needed to train LLMs. P15, reflecting that they may be one of one million authors, noted that “I don’t actually think that the companies who have created these models own me a large amount of money” even if the amount of money needed to compensate all authors would be significant. Other writers noted that they wouldn’t even necessarily want a small amount of compensation: “If it was gonna be 23 cents, because that’s really how much it’s worth, in some sense like… I think I should care more. But I don’t actually want a check for 23 cents” (P7). P32 noted that a very small amount of compensation in fact shouldn’t be considered compensation at all—compensation necessitates that it compensates the writer, and a small amount wouldn’t actually be compensation for their work.
@@
I guess, because no one will go look at it. They won’t. They won’t care. If I’m going to write something, I just want to use the model. I just want to chat, write in my prompt and get my results. I’m not going to take the time to go, Gee! I wonder what 864,000 people were used to train this model. (P24)
A few writers did consider ways to deal with the scale of LLM training data. P18 wanted credit to be based on the output, such that outputs that were ”more influenced” by some writing could highlight the particularly relevant contributions. P15 suggested that rather than consider compensation at the individual level, “the moral thing to do would be to figure out some way to offset the disruption you’re causing. The way that you would do that isn’t necessarily just by paying everybody whose work has been fed into you know… it would be some other grant or taxation.” Similarly P33 noted that industry disruption felt more salient than disruption to their individual career. P14 suggested that something like stock options would allow them to be compensated in line with the value that was created.

### v2

@@ 5.1.2. Respect
While writers understood that by publishing their work they are allowing others to use that work in perhaps unexpected ways, they also felt that the use of their work as training data indicated a lack of respect both for writing as an endeavor and writers as people. Although writers did not express that all uses of their writing require consent—for instance, as in the section above about the creative chain, they expect that others will learn from their work or perhaps even try to mimic it—the use of their writing as training data felt different. P6, a professional romance novelist, said that the use of their writing for AI may be more like signing away their foreign subrights, which would require consent and be part of contract negotiations when publishing.
Mechanisms of respect: compensation is key for professionalized writers. The idea that writers wanted respect was very much tied to the mechanisms in which respect can be shown. The mechanisms that writers desired was a function of many things: their degree of professionalization, the type of writing they worked on, their sense of precarity in the labor market, and their perception of the organizations that were or could be using their writing. Primarily, writers wanted respect for writing as a laborious act on its own—to write a book takes time, effort, and expertise. As P9 (professional romance novelist) put it,
@@
The people we talked to who were professional writers—who earned the majority of the income from their creative writing practice—were most likely to cite compensation as the main mechanism of respect they would expect. For instance, P8, a professional freelance writer, noted that, “And I say this as a, you know, a very pro AI person who thinks it’s the next creative way. But content producers actually need to be compensated on some level.” P11 (professional journalist and novelist) made the comparison with other kinds of expertise: “If NASA wanted my expertise to hire me to help in the space program they would actually have to pay me quite a lot of money, and they would do so.” However, technology companies were not treating writers as if they brought any kind of expertise or as if their labor was valuable.
@@
Power differentials modulate desire for compensation. Writers were also quick to point out the large power differential between the corporations creating these models and the writers trying to earn a living. For instance, P7 (professional journalist) explains that, “The tech companies are making money hand over fist, and I feel like they can compensate people.” Professional writers were likely to say they would require compensation regardless of whether or not the model creator was making money, but many writers did make a distinction between organizations they perceived could afford to compensate writers, and non-profit organizations or applications. We return to the idea of who can afford to compensate writers in the realistic expectations section.
@@
Different kinds of writing require different kinds of respect. Writers discussed genre distinctions in terms of the level of labor involved. P9, a professional romance novelist, talked about the use of certain kinds of books feeling particularly disturbing, “like Pulitzer Prize-winning work that’s happening in nonfiction,” which represents decades of research then being taken without consent or compensation and made “part of the churn.” In this case, writers wanted respect for time-consuming labor that writing represents.
@@
Other writers noted that the use of certain kinds of writing felt more disrespectful than others because of the contents of the writing, rather than the labor taken to produce it. P4 (fan fiction writer and poet) talked about some writing being “deeply personal” and being “disturbed by the lack of care” that is shown by the act of putting such work into a language model. P32, in discussing a colleague who had written a nonfiction book which included accounts of assault, noted that their feelings and words about the person who assaulted them now lives on in the predictive models trained on that book, saying, ”There is an aspect of that that is like literal haunting.” Here, respect was not about compensation but about understanding the sensitive nature of some kinds of writing.
@@
How to value writing. Writers discussed how they might consider whether or not to contribute certain works of theirs, which led them to express how different kinds of writing may be valued differently. P14 (poet and translator) said that if they were asked for consent, “there are things that I would probably hold back in order to wait until I’ve gotten the value out of them and then at that point I wouldn’t worry about it as much.” Similarly P13 (fan fiction and science fiction writer) noted that they used to post a lot more of their ideas on social media before they started winning awards for their stories; now, they consider their ideas to be more valuable and are more hesitant about sharing them publicly. Other writers noted that older writing is often less valuable, either, as P14 noted, because they have already gotten value out of it, or because older writing is worse—they have become a better writer and therefore value their more recent writing more (P7, journalist and novelist).
@@ 5.1.3. The Human Element
Many writers see the text generation of LLMs as an attempt to mimic human writing, and, in many cases, an author’s unique style or voice. This is tied to discussions of the creative chain (subsubsection 5.1.1) and writers view the training of LLMs on a particular writer’s voice as impoverished and decontextualized. When they read writing that inspires them, writers appreciate, contextualize, question, compare, and transform a text within their own social, embodied, and emotional contexts, something mass aggregates of textual data cannot do. P6 (professional /novelist) contrasts these processes: “My absolute favorite book was instrumental in the way that I wrote the favorite of all of the books I’ve written. But it’s still coming through me. Whereas, I feel like ChatGPT would end up poorly mimicking me, as opposed to being something that came from someone else, but was formed by their appreciation for me.”
@@
Implicit in these distinctions is an understanding that LLMs do not have authorial intent, which is central to writing as a human activity. That LLMs do not have authorial intent when generating text is viewed as an asset to those writers choosing to use them in creative ways. P31, who has used LLMs in their creative practice, notes this explicitly, while problematizing how LLMs are colloquially talked about as “collaborators” in some contexts. They say:
@@
The audience for a training dataset of creative work is not the audience that writers want or imagine. Undergirding our conversations with writers was the understanding that the imagined or intended audience for a writer’s work is other people, not a dataset using writing to train an LLM. Every direct mention of the words ‘audience’ or ‘reader’ in our interviews was in reference to an explicitly human audience. At least one of these two terms was mentioned in all of the interviews we conducted.
In light of the strong connection between humanity and writers’ conceptions of audience, writers were left to conjecture about who the ‘audience’ of training data really was. Algorithms, corporations, and users of LLMs were all variously floated as ideas of what might constitute the readership for training datasets. P26 wants to share their work with people, but is “not ready to put [even] a little piece of it into the algorithm.” The scale and combinatory nature of training datasets provoked skepticism and sometimes fear around this imagined ‘readership,’ quite unlike the readerships writers imagine for themselves.
@@ 5.2.1. Lack of Control
Certain uses of writing are restricted and there are avenues for recourse. Writers often sign contracts that outline a publisher’s right to republish a book in a foreign country, or create a movie adaption, indicating that writers do have control over a variety of use cases of their writing. For instance, P11 (a professional novelist and journalist) said, “There are lots of sources of pirated books on the Internet, but they’re illegal, and we try to shut them down. If ever I find a free copy of my book circulated somewhere, I’ll ask my publisher to send a cease and desist letter, and it usually gets taken down.
@@
It is difficult to understand how to restrict only certain use cases of LLMs. Writers did not always disagree with the creation of language models per se, and instead had concerns only about certain use cases. But many writers noted that it seems impossible to prevent certain use cases. For instance, P24 (a fantasy novelist) would consider consenting to the use of their writing if they could ensure that the model would not be used to compete with them, but felt “like that would be an impossible ask. People are just gonna lie and go ‘No, I’m not writing a novel.’” Others noted that while it may be possible to audit big companies, “there are all these people putting out small models, anyone can do this now, and it almost feels impossible to rein in” (P2, a writer experienced with LLMs).
@@
This is what tech companies have always done. Uber is like, we’re just gonna make these cars. And we’ll see if cities can catch up with us. Or Google Maps when they’re like, we’re just gonna photograph everybody’s house. And then we’ll see if anyone gets mad. I think that’s totally what AI writing software is doing. They’re waiting to see what people are gonna do. (P12)
@@
Finally, writers acknowledged the difficulty of predicting future use cases. As technology changes, it can be hard to know what you might be agreeing to in the moment. Some established professional novelists were especially concerned about those at the start of their careers—established writers were concerned about up-and-coming authors (P9, P12, P20) and writers we interviewed who were at the start of their careers noted their particular predicament (P13). As an early stage writer, one has less power to negotiate certain terms out of their contract and may end up agreeing to terms that, years down the line, perhaps when the writer is more famous, are unexpected, unprecedented, and ultimately undesirable.
5.2.2. Industry Impacts
@@
Though public discourse about the impact of language models on the creative writing industry often takes place in the future, writers described present day impacts of language models on their careers and fields. P12 described the strain language models have placed on literary editors, who “spend hours every week” sifting through large amounts of “junk pitches and generated stories.” P12 recalled how this influx of generated work led one prominent literary magazine (Clarkesworld) to “close submissions entirely,” which “really hurts people who are just getting started, people who already have this narrowing slate of options where they can submit their work.” Other writers expressed similar concern over the impacts of language models on new writers getting started in the industry (P9).
@@
Writers with academic teaching positions described confronting the impacts of language models on creative writing education. These writers explained how their creative writing careers and teaching careers are inextricable, either because they are hired to teach because of their creative writing careers, or because they see teaching as part of their creative practice. These writers note that even if they are morally against LLMs, for instance because of the energy used to train and use them, they must deal with this technology in their classrooms (P12, P14, P16, P18, P32).
@@
Some writers also described positive impacts of language models and AI on their careers. Self-published authors described using language models and other forms of generative AI to help with the work a traditional publisher might do, like creating a cover or producing marketing materials for their work (P24, P27). P31, who makes use of language models in their work, described how their career has benefited from advances in LLMs and the resulting increase in public interest in AI.
@@
Predicted future impacts. By and large, most participants found that in their current state, language models are not direct competition for writers as they do not produce sufficiently sophisticated, high quality writing. When speculating about how future advanced versions of language models might impact their careers and industries, many writers imagined competing with a language model that had been trained to mimic their writing. Other writers speculated more broadly about how language models might compete with or displace writers, i.e., ”an AI that writes better novels than humans” (P12, professional novelist).
@@
I don’t think [LLMs are] particularly useful for the creative industry. I mean, I could see them being useful for, like, … discovering a new kind of a new antibiotic or something like that… But, human expression? I guess I think of poetry or fiction as a deeply human activity. And it sort of baffles me why people would want to read what has been historically a human activity that has been produced by a machine. (P10, poet and essayist)
@@
Backdrop of a changing industry. Some writers acknowledged that their resistance to contribute to training datasets was “compounded by a sense of precarity” in the creative writing industry (P3, poet). P20 (professional science fiction novelist) noted that “it’s a really difficult time for people in our profession and for creative industries in general. The job market is constantly shrinking.
Some writers expressed protectiveness towards the creative industry, citing the value of humanity intrinsic to creative work. The questions these writers were asking were existential, implying that the existence of AI-generated art and literature is forcing us to ask, ‘What kind of world do we want to live in?’ P9 (professional romance novelist) notes:
@@
P13, along with several other writers we spoke with, wants to live in a world where people can make a living as writers: “I don’t think it’s good to make it cheaper to make art for the fact that people need to be employed and I would prefer to be employed as a writer than as a lawyer.
@@ 5.2.3. Interpretation of Scale
Compensation is hard to determine and likely to be insignificant given the scale of LLMs. While most writers objected to their data being collected by corporations, they simultaneously had difficulty comprehending how to value their contribution when considering the scale of training data used. Some noted that their style was not distinct enough to be noticeable in any outputs or to significantly change the model (P2, P24). P10 (poet and essayist) described this as, “It’s hard for me to trace how money being made is based on the training data.” Others felt it was hard to feel personally slighted when their contribution is so very small (P29, fan fiction writer). Still others were unsure exactly how compensation would even work: “It’s not like a Spotify thing where you can be like, Okay, they listen to this track, or they read this chapter” (P17).
Despite many writers wanting compensation for the use of their work, they acknowledged that such compensation is likely to be small given the amount of data needed to train LLMs. P15, who has worked with LLMs in their creative practice, reflecting that they may be one of one million authors, noted that “I don’t actually think that the companies who have created these models owe me a large amount of money” even if the amount of money needed to compensate all authors would be significant. Other writers noted that they wouldn’t even necessarily want a small amount of compensation: “If it was gonna be 23 cents, because that’s really how much it’s worth, in some sense like… I think I should care more. But I don’t actually want a check for 23 cents” (P7). P32 noted that a very small amount of compensation in fact shouldn’t be considered compensation at all—compensation necessitates that it compensates the writer, and a small amount wouldn’t actually be compensation for their work.
@@
I guess, because no one will go look at [a list of contributing authors]. They won’t. They won’t care. If I’m going to write something, I just want to use the model. I just want to chat, write in my prompt and get my results. I’m not going to take the time to go, Gee! I wonder what 864,000 people were used to train this model. (P24)
A few writers did consider ways to deal with the scale of LLM training data. P18 (fan fiction writer) wanted credit to be based on the output, such that outputs that were ”more influenced” by some writing could highlight the particularly relevant contributions. P15 suggested that rather than consider compensation at the individual level, “the moral thing to do would be to figure out some way to offset the disruption you’re causing. The way that you would do that isn’t necessarily just by paying everybody whose work has been fed into you know… it would be some other grant or taxation.” Similarly P33 noted that industry disruption felt more salient than disruption to their individual career. P14 suggested that something like stock options would allow them to be compensated in line with the value that was created.